# Doubling the size of quantum selected configuration interaction based on seniority-zero space and its application to QC-QSCI-AFQMC

### v1

@@ I Introduction
Quantum chemistry on quantum computers has advanced rapidly over the past decade [Peruzzo2014variational, Kandala2017hardware, Huggins2022unbiasing, Kanno2023quantum, Robledo-Moreno2025chemistry]. While the rapid development of quantum hardware has been crucial to this progress, improvements in quantum algorithms—particularly refinements to quantum–classical hybrid workflows—have been equally important. Although current quantum-chemical calculations on quantum computers are typically approximate, they are now reaching problem sizes for which classically exact solutions are no longer feasible on conventional (classical) computers. They are beginning to compete with state-of-the-art classical methods, such as coupled-cluster (CC) and density matrix renormalization group (DMRG) calculations. In light of these developments, performing practically relevant calculations is becoming realistic on today’s quantum hardware.
@@
Despite this success, further improvements are still needed to make more effective use of QSCI [Nakagawa2024ADAPT, Sugisaki2025Hamiltonian, Mikkelsen2024Quantum, Chen2026neural]. In particular, the standard fermion-to-qubit mapping, in which the number of qubits scales one-to-one with the number of spin-orbitals, drastically increases the size of the many-qubit Hilbert space and the number of possible entangling operations, pushing circuit depths well beyond what is feasible. Constant-factor reductions achieved through symmetry-based qubit tapering [Bravyi2017tapering, Seria2020reducing] may not be sufficient. Aggressive compression that disregards the structure and symmetries of Hilbert space risks undermining the design principles of quantum circuits.
@@
A natural way to partition Hilbert space is by the seniority number, a long-standing concept in quantum chemistry [Bytautas2011seniority, Bytautas2015seniority] as well as in nuclear and condensed-matter physics. The seniority number is defined as the number of unpaired electrons in a Slater determinant, and the seniority-zero subspace consists of electronic configurations with no unpaired electrons. By retaining only paired electrons, the seniority-zero space captures the pair correlations essential for static correlation and is known to describe static correlation effectively. Moreover, doubly occupied configuration interaction (DOCI) [Allen1962electron, Smith1965natural, Weinhold1967reduced, Veillard1967complete, Cook1975doubly, Fantucci1977direct, Couty1997generalized, Kollmar2003new], i.e., configuration interaction (CI) restricted to the seniority-zero space, involves far fewer configurations than full CI, making it more tractable as the number of orbitals increases.
@@
On the other hand, it is known that the seniority-zero space captures only a limited subset of electronic correlation. Achieving fully quantitative accuracy ultimately requires enabling seniority-breaking, and the seniority-zero space provides a practical foundation for this.
@@
The feasibility of utilizing the seniority-zero space for quantum computation has been explored [Elfving2011simulating, Kottmann2022optimized, Kim2024variational, Zhao2023orbital, Zhao2024enhancing, O’Brien2023purification, Khamoshi2021correlating, Khamoshi2023AGP, Khamoshi2023state]. The primary advantage is that the number of qubits can be made equal to the number of spatial orbitals rather than the number of spin orbitals, effectively halving the qubit count. Consequently, this also enables compact and readily implementable quantum-circuit designs. Variational algorithms, circuit implementations, and state preparation for QPE have been studied. Therefore, utilizing the seniority-zero space in quantum computing is a promising research direction, offering avenues for further algorithmic improvements.
@@
In this paper, we propose a QSCI method that samples from a seniority-zero space. Moreover, we incorporate classical post-processing via a ph-AFQMC calculation to elevate QSCI to a level of highly accurate quantum-chemical calculations. As one approach to improving accuracy through seniority breaking, we expand each sampled result into two spin-resolved configurations and treat them in a Cartesian product space. The resulting QSCI wave function is used as the trial wave function for the ph-AFQMC calculation to recover electronic correlation—primarily dynamical correlation—across the full orbital space. The proposed method is tested on benchmark molecular systems, including a hydrogen chain and N2. The present method is also applied to the oxygen-addition reaction of a BODIPY dye, demonstrating its applicability to a practical molecular system.
@@ II Computational methods
II.1 Quantum selected configuration interaction and its extension
@@ II.1.1 Basic formalism of QSCI
In this section, we provide a brief review of the basic formalism of QSCI [Kanno2023quantum]. The QSCI method constructs the effective Hamiltonian {align} ^H_eff = ^P_S^H^P_S and then obtains its eigenstates. Here, $\hat{H}$ is the electronic structure Hamiltonian, and in the case of active-space QSCI, it corresponds to the active-space Hamiltonian. This effective Hamiltonian is constructed based on the sampling results from a quantum computer. It defines a subspace $S$ spanned by the computational-basis states associated with the quantum state of interest. The operator $\hat{P}_{S}$ is the projector onto $S$, and can be expressed as {align} ^P_S = ∑_i=1^R | Φ_i ⟩⟨Φ_i | using the computational basis states $\{|\Phi_{i}\rangle\}$ spanning the subspace $S$.
The eigenstates of the effective Hamiltonian $\hat{H}_{\text{eff}}$ are obtained by solving {align} ^H_eff | Ψ⟩= E | Ψ⟩ on a classical computer, where $|\Psi\rangle$ and $E$ are the eigenvector and the corresponding eigenvalue, respectively.
@@ II.1.3 DOCI-QSCI
Thus, although the sampling is performed in the DOCI (seniority-zero) representation, the subspace used to construct the effective Hamiltonian in QSCI is not restricted to seniority-zero. This procedure reduces the cost of determinant selection by sampling in an $N$-orbital representation, while the effective Hamiltonian is constructed in the Cartesian product space guided by the seniority-zero sampling.
@@ III Computational details
Self-consistent field (SCF) calculations for generating molecular integrals, as well as conventional quantum-chemical calculations including coupled-cluster singles and doubles (CCSD) and CCSD with perturbative triples (CCSD(T)), were performed using PySCF [PySCF, PySCF2]. We used the cc-pVTZ, cc-pVQZ, and 6-31G(d,p) basis-sets for H6 chain, N2, and BODIPY–O2 complex, respectively. Reference heat-bath configuration interaction (HCI) calculations were performed using PyCI [PyCI].
Quantum computations were carried out using the Qiskit library v2.2.3 [qiskit2024]. We employed the spinless unitary cluster Jastrow (UCJ) ansatz with a single repetition. The Jastrow interactions were restricted to adjacent orbitals (qubits), i.e., in a local unitary cluster Jastrow (LUCJ) manner [Motta2023bridging]. The circuit parameters were fixed to the active-space amplitudes obtained from frozen-core CCSD calculations. We used the IBM Quantum device ibm_kobe and the noiseless simulator. When using the real quantum device, sampling results that do not preserve the number of electrons are physically meaningless and are discarded.
@@
The ph-AFQMC calculations were conducted using ipie v0.6.2 [Malone2023ipie]. We employed a time step of 0.005 $E_{h}^{-1}$, 50 steps, and 3000 blocks unless otherwise specified.
The reaction path for the addition of singlet oxygen to the BODIPY dye was investigated at the RB3LYP/6-31G(d) level and identified using the intrinsic reaction coordinate calculations. Density functional theory calculations were carried out using Gaussian 16 Rev. C [g16].
@@ IV Results and discussion
Here, we present the computational results of DOCI-QSCI-AFQMC and discuss their validity. First, we present results for the dissociation of the H6 chain and N2, typical molecular systems for strong electron correlation. Next, we examine the cyclization reaction of a BODIPY molecule with singlet oxygen as a showcase of a realistic chemical reaction.
@@ IV.1 H6 chain
Figure 2(a) and (b) show the energy curves and the energy differences relative to HCI for the one-dimensional H6 chain, including DOCI-QSCI-AFQMC results obtained with a classical simulator and on the quantum computer ibm_kobe. For DOCI-QSCI-AFQMC calculations performed with the simulator, we also report results in which the sampled subspace is further enlarged using a conventional selected-CI procedure. We used 400 walkers in the ph-AFQMC calculations.
@@
In contrast, DOCI-QSCI-AFQMC using a noiseless simulator does not reproduce the high-accuracy results unless the subspace is enlarged in the selected CI manner. This reflects the fact that, with a finite number of samples, the probability distribution defined by the ansatz can be heavily skewed toward a small number of bitstrings, such as the Hartree-Fock configuration, leading to insufficient coverage of the relevant determinants. However, highly accurate energy values can be readily obtained by expanding the space.
The control calculation using a DOCI trial wave function (DOCI-AFQMC) indicates that DOCI alone does not yield quantitatively reliable results for this strongly correlated system, consistent with a previous study [Yoshida2025auxiliary]. DOCI-AFQMC yields a larger energy error than DOCI-QSCI-AFQMC, beyond the statistical uncertainty of the quantum Monte Carlo sampling. This highlights the difference between the seniority-zero space and the Cartesian product space of bitstrings guided by seniority-zero sampling.
@@ IV.2 Potential energy curve for N2
For DOCI-QSCI-AFQMC and DOCI-QSCI, we present two variants: (i) a fixed subspace constructed from sampling (Fixed), and (ii) an enlarged subspace obtained by further expanding the space using a conventional selected-CI procedure (Enlarged).
The $10^{5}$ shots for each point are insufficient; consequently, DOCI-QSCI with the fixed subspace does not reproduce the dissociation curve quantitatively. Accordingly, DOCI-QSCI-AFQMC, which uses the corresponding DOCI-QSCI trial wave function, also fails to describe the dissociation accurately. This indicates that the ph-AFQMC post-processing inherits the deficiency in static correlation present in the trial state.
@@
By contrast, RCCSD and RCCSD(T) yield qualitatively incorrect energy curves. This well-known breakdown of single-reference methods underscores the need for multireference treatments.
@@ IV.3 Reaction between BODIPY dye and singlet oxygen
In this section, we apply the present method to the addition of singlet oxygen to a BODIPY molecule. The addition of singlet oxygen to conjugated systems has attracted interest in modern high-level quantum chemical calculations [Winslow2024spin]. To describe configuration interactions associated with bond formation and cleavage, as well as the stabilization of conjugated systems, more orbitals must be included in the active space, thereby increasing computational cost. BODIPY dyes are among the most prominent conjugated systems in modern chemistry, and understanding their decomposition pathways is crucial for extending their lifespan [Mula2008design, Jagtap2013change, Rybczynski2021photochemical]. Here, we explore the feasibility of treating large active spaces using quantum computing.
@@
The activation and reaction energies $(E_{a},E_{r})$ are shown in Figure 4 and Table 1. Density fitting was used for calculations other than RB3LYP. DOCI-QSCI and DOCI-QSCI-AFQMC employ natural orbitals obtained from a CISD calculation; therefore, increasing the active space size is expected to improve the accuracy of the resulting energies systematically. The number of shots for sampling is $10^{6}$. For the ph-AFQMC calculations, we used 512 walkers and 1000 blocks.
@@
The DOCI-QSCI-AFQMC results show less dependence on the active space size than the corresponding DOCI-QSCI results. This suggests that performing ph-AFQMC in the full orbital space reduces the sensitivity of the computed energies to the choice of active space. The resulting activation energy is predicted to be similar to RB3LYP.
@@ V Conclusions
We proposed DOCI-QSCI, an improved variant of QSCI obtained by sampling in the seniority zero space. We also proposed using the DOCI-QSCI wave function as the trial wave function in the QC-QSCI-AFQMC framework [Yoshida2025QSCIAFQMC, Danilov2025enhancing]; we refer to this approach as DOCI-QSCI-AFQMC. Working in the seniority-zero space allows the number of qubits to be equal to the number of spatial orbitals rather than the number of spin orbitals, thereby halving the required number of qubits (or, equivalently, doubling the number of spatial orbitals that can be treated for a fixed qubit budget). On the other hand, because the seniority-zero space is smaller than the full Hilbert space, restricting the wave function to this subspace can affect quantitative accuracy. In this paper, we considered Cartesian-product states constructed from the sampled bitstrings. We evaluated the QSCI energy in the resulting product space, which extends beyond the corresponding seniority-zero subspace.
The proposed methods were tested on several molecular systems. In the H6 chain benchmark, DOCI-QSCI-AFQMC provided accurate results both with real quantum hardware and in simulation. In some simulations, we used bitstrings obtained from low-shot sampling as seeds and enlarged the space by adding $\alpha/\beta$ strings in a heat-bath CI fashion. We also confirmed that the DOCI-QSCI trial wave function was of higher quality than the seniority-zero DOCI trial wave function.
@@
Overall, DOCI-QSCI and DOCI-QSCI-AFQMC, which reduce the required number of qubits, are expected to facilitate the scaling up of practical quantum chemical calculations using quantum computers. On the other hand, open-shell and high-spin states are not accessible within the seniority-zero space, and further methodological development is required.

### v2

# arXiv:2602.07912v2[quant-ph] 27 Feb 2026

@@ I Introduction
Quantum chemistry on quantum computers has advanced rapidly over the past decade [Peruzzo2014variational, Kandala2017hardware, Huggins2022unbiasing, Kanno2023quantum, Robledo-Moreno2025chemistry]. While the rapid development of quantum hardware has been crucial to this progress, improvements in quantum algorithms—particularly refinements to quantum–classical hybrid workflows—have been equally important. Although current quantum-chemical calculations on quantum computers are typically approximate, they are now approaching problem sizes for which finding classically exact solutions is no longer feasible on conventional (classical) computers. They are beginning to compete with state-of-the-art classical methods, such as coupled cluster (CC) and density matrix renormalization group calculations. In light of these developments, performing practically relevant calculations is becoming realistic on today’s quantum hardware.
@@
Despite this success, further improvements are still needed to make more effective use of QSCI [Nakagawa2024ADAPT, Sugisaki2025Hamiltonian, Mikkelsen2024Quantum, Chen2026neural]. In particular, the standard fermion-to-qubit mapping, in which the number of qubits scales one-to-one with the number of spin orbitals, drastically increases the size of the many-qubit Hilbert space and the number of possible entangling operations, pushing circuit depths well beyond what is feasible. Constant-factor reductions achieved through symmetry-based qubit tapering [Bravyi2017tapering, Seria2020reducing] may not be sufficient. Aggressive compression that disregards the structure and symmetries of the Hilbert space risks undermining the design principles of quantum circuits.
@@
A natural way to partition Hilbert space is by the seniority number, a long-standing concept in quantum chemistry [Bytautas2011seniority, Bytautas2015seniority] and in nuclear and condensed-matter physics. The seniority number is defined as the number of unpaired electrons in a Slater determinant, and the seniority-zero subspace comprises electronic configurations with no unpaired electrons. By retaining only paired electrons, the seniority-zero space captures the pair correlations essential for static correlation and is known to describe static correlation effectively. Moreover, doubly occupied configuration interaction (DOCI) [Allen1962electron, Smith1965natural, Weinhold1967reduced, Veillard1967complete, Cook1975doubly, Fantucci1977direct, Couty1997generalized, Kollmar2003new]—i.e., configuration interaction (CI) restricted to the seniority-zero space—involves far fewer configurations than full CI, making it more tractable as the number of orbitals increases.
@@
Conversely, it is known that the seniority-zero space captures only a limited subset of electronic correlation. Achieving fully quantitative accuracy ultimately requires enabling seniority breaking, and the seniority-zero space provides a practical foundation for this.
@@
The feasibility of utilizing the seniority-zero space for quantum computation has been explored [Elfving2011simulating, Kottmann2022optimized, Kim2024variational, Zhao2023orbital, Zhao2024enhancing, O’Brien2023purification, Khamoshi2021correlating, Khamoshi2023AGP, Khamoshi2023state]. The primary advantage is that the number of qubits can be equal to the number of spatial orbitals rather than the number of spin orbitals, effectively halving the qubit count. Consequently, this also enables compact and readily implementable quantum-circuit designs. Variational algorithms, circuit implementations, and state preparation for QPE have been studied. Therefore, utilizing the seniority-zero space in quantum computing is a promising research direction, offering avenues for further algorithmic improvements.
@@
In this paper, we propose a QSCI method that samples from the seniority-zero space. Moreover, we incorporate classical postprocessing via a ph-AFQMC calculation to elevate QSCI to highly accurate quantum-chemical calculations. As one approach to improving accuracy through seniority breaking, we expand each sampled result into two spin-resolved configurations and treat them in a Cartesian product space. The resulting QSCI wave function is used as the trial wave function for the ph-AFQMC calculation to recover electronic correlation—primarily dynamical correlation—across the full orbital space. The proposed method is tested on benchmark molecular systems, such as a hydrogen chain and N2. The present method is also applied to the oxygen addition reaction of a BODIPY dye, demonstrating its applicability to a practical molecular system.
@@ II Computational methods
II.1 Quantum-selected configuration interaction and its extension
@@ II.1.1 Basic formalism of QSCI
This section provides a brief review of the basic formalism of QSCI [Kanno2023quantum]. The QSCI method constructs the effective Hamiltonian {align} ^H_eff = ^P_S^H^P_S and then obtains its eigenstates. Here, $\hat{H}$ is the electronic structure Hamiltonian, and in the case of active-space QSCI, it corresponds to the active-space Hamiltonian. This effective Hamiltonian is constructed based on the sampling results from a quantum computer. It defines a subspace $S$ spanned by the computational basis states associated with the quantum state of interest. The operator $\hat{P}_{S}$ is the projector onto $S$, and can be expressed as {align} ^P_S = ∑_i=1^R | Φ_i ⟩⟨Φ_i | using the computational basis states $\{|\Phi_{i}\rangle\}$ spanning the subspace $S$.
The eigenstates of the effective Hamiltonian $\hat{H}_{\text{eff}}$ are obtained by solving the following equation: {align} ^H_eff | Ψ⟩= E | Ψ⟩ on a classical computer, where $|\Psi\rangle$ and $E$ are the eigenvector and the corresponding eigenvalue, respectively.
@@ II.1.3 DOCI-QSCI
Thus, although the sampling is performed in the DOCI (seniority-zero) representation, the subspace used to construct the effective Hamiltonian in QSCI is not restricted to seniority zero. This procedure reduces the cost of determinant selection by sampling in an $N$-orbital representation, while the effective Hamiltonian is constructed in the Cartesian product space guided by the seniority-zero sampling.
@@ III Computational details
Self-consistent field (SCF) calculations for generating molecular integrals, as well as conventional quantum-chemical calculations including CC singles and doubles (CCSD) and CCSD with perturbative triples (CCSD(T)), were performed using PySCF [PySCF, PySCF2]. We used the cc-pVTZ, cc-pVQZ, and 6-31G(d,p) basis sets for the H6 chain, N2, and BODIPY–O2 complex, respectively. Reference heat-bath configuration interaction (HCI) calculations were performed using PyCI [PyCI].
Quantum computations were performed using the Qiskit library v2.2.3 [qiskit2024]. We employed the spinless unitary cluster Jastrow (UCJ) ansatz with a single repetition. The Jastrow interactions were restricted to adjacent orbitals (qubits); i.e., in a local UCJ manner [Motta2023bridging]. The circuit parameters were fixed to the active-space amplitudes obtained from frozen-core CCSD calculations. We used an IBM Quantum device ibm_kobe and a noiseless simulator. When using the real quantum device, sampling results that do not preserve the number of electrons are physically meaningless and are discarded.
@@
The ph-AFQMC calculations were conducted using ipie v0.6.2 [Malone2023ipie]. We employed a time step of 0.005 $E_{h}^{-1}$, 50 steps, and 3000 blocks, unless otherwise specified.
The reaction path for the addition of singlet oxygen to the BODIPY dye was investigated at the RB3LYP/6-31G(d) level and identified by intrinsic reaction coordinate calculations. Density functional theory calculations were carried out using Gaussian 16 Rev. C [g16].
@@ IV Results and discussion
Here, we present the computational results of DOCI-QSCI-AFQMC and discuss their validity. First, we present results for the dissociation of the H6 chain and N2, typical molecular systems for strong electron correlation. Next, we examine the cyclization reaction of a BODIPY molecule with singlet oxygen as an example of a realistic chemical reaction.
@@ IV.1 H6 chain
Figure 2(a) and (b) show the energy curves and the energy differences relative to HCI for the one-dimensional H6 chain, including DOCI-QSCI-AFQMC results obtained with a classical simulator and on the quantum computer ibm_kobe. For DOCI-QSCI-AFQMC calculations used by the simulator, we also report results in which the sampled subspace is further enlarged using a conventional selected-CI procedure. We used 400 walkers in the ph-AFQMC calculations.
@@
In contrast, DOCI-QSCI-AFQMC using a noiseless simulator does not reproduce the high-accuracy results unless the subspace is enlarged in the selected-CI manner. This reflects the fact that with a finite number of samples, the probability distribution defined by the ansatz can be heavily skewed toward a small number of bitstrings, such as the Hartree-Fock configuration, leading to insufficient coverage of the relevant determinants. However, highly accurate energy values can be readily obtained by expanding the space.
The control calculation using a DOCI trial wave function (DOCI-AFQMC) indicates that DOCI alone does not provide quantitatively reliable results for this kind of strongly correlated system, which is consistent with a previous study [Yoshida2025auxiliary]. DOCI-AFQMC yields a larger energy error than DOCI-QSCI-AFQMC, beyond the statistical uncertainty of quantum Monte Carlo sampling. This highlights the difference between the seniority-zero space and the Cartesian product space of bitstrings guided by seniority-zero sampling.
@@ IV.2 Potential energy curve for N2
For DOCI-QSCI-AFQMC and DOCI-QSCI, we present two variants: (i) a fixed subspace constructed from sampling (Fixed) and (ii) an enlarged subspace obtained by further expanding the space using a conventional selected-CI procedure (Enlarged).
The $10^{5}$ shots for each point are insufficient; consequently, DOCI-QSCI with a fixed subspace does not reproduce the dissociation curve quantitatively. Similarly, DOCI-QSCI-AFQMC, which uses the corresponding DOCI-QSCI trial wave function, also fails to describe the dissociation accurately. This indicates that the ph-AFQMC postprocessing inherits the deficiency in static correlation present in the trial state.
@@
In contrast, RCCSD and RCCSD(T) yield qualitatively incorrect energy curves. This well-known breakdown of single-reference methods underscores the need for multireference treatments.
@@ IV.3 Reaction between BODIPY dye and singlet oxygen
In this section, we apply the presented method to the addition of singlet oxygen to a BODIPY molecule. The addition of singlet oxygen to conjugated systems has attracted interest in modern high-level quantum chemical calculations [Winslow2024spin]. To describe configuration interactions associated with bond formation and cleavage, as well as the stabilization of conjugated systems, more orbitals must be included in the active space, which increases the computational cost. BODIPY dyes are among the most prominent conjugated systems in modern chemistry, and understanding their decomposition pathways is crucial for extending their lifespan [Mula2008design, Jagtap2013change, Rybczynski2021photochemical]. Here, we explore the feasibility of treating large active spaces using quantum computing.
@@
The activation and reaction energies $(E_{a},E_{r})$ are shown in Figure 4 and Table 1. Density fitting was used for calculations other than RB3LYP. DOCI-QSCI and DOCI-QSCI-AFQMC employ natural orbitals obtained from a CISD calculation; increasing the size of active space is therefore expected to improve the accuracy of the resulting energies. The number of shots for sampling is $10^{6}$. For the ph-AFQMC calculations, we used 512 walkers and 1000 blocks.
@@
The DOCI-QSCI-AFQMC results show less dependence than the corresponding DOCI-QSCI results on the size of the active space. This suggests that performing ph-AFQMC in the full orbital space reduces the sensitivity of the computed energies to the choice of active space. The resulting activation energy is predicted to be similar to RB3LYP.
@@ V Conclusions
We have proposed DOCI-QSCI, an improved variant of QSCI obtained by sampling in the seniority-zero space. We have also proposed using the DOCI-QSCI wave function as the trial wave function in the QC-QSCI-AFQMC framework [Yoshida2025QSCIAFQMC, Danilov2025enhancing]; we refer to this approach as DOCI-QSCI-AFQMC. Working in the seniority-zero space allows the number of qubits to be equal to the number of spatial orbitals rather than the number of spin orbitals, thereby halving the required number of qubits (or, equivalently, doubling the number of spatial orbitals that can be treated for a fixed qubit budget). Conversely, because the seniority-zero space is smaller than the full Hilbert space, restricting the wave function to this subspace can affect quantitative accuracy. In this paper, we considered Cartesian-product states constructed from the sampled bitstrings. We evaluated the QSCI energy in the resulting product space, which extends beyond the corresponding seniority-zero subspace.
The proposed methods were tested on several molecular systems. In the H6 chain benchmark, DOCI-QSCI-AFQMC provided accurate results with real quantum hardware and in simulation. In some simulations, we used bitstrings obtained from low-shot sampling as seeds and enlarged the space by adding $\alpha/\beta$ strings in a heat-bath CI fashion. We also confirmed that the DOCI-QSCI trial wave function was of higher quality than the seniority-zero DOCI trial wave function.
@@
Overall, DOCI-QSCI and DOCI-QSCI-AFQMC, which reduce the required number of qubits, are expected to facilitate the scaling up of practical quantum chemical calculations using quantum computers. Conversely, open-shell and high-spin states are not accessible within the seniority-zero space, and further methodological development is required.